OpenStudy (susanna):

State the period and phase shift of the function y=3 tan (4x - pi/3).

OpenStudy (susanna):

@SolomonZelman

OpenStudy (michele_laino):

hint: we can rewrite such function as below: \[\huge y = 3\tan \left( {4\left( {x - \frac{\pi }{{12}}} \right)} \right)\]

OpenStudy (michele_laino):

now, you can read the period and the phase shift

OpenStudy (susanna):

period is pi/4 and phase shift is -pi/12 ??

OpenStudy (michele_laino):

that's right! Better is to write the phase shift without the minus sign

OpenStudy (susanna):

so is the phase shift positive or negative?

OpenStudy (michele_laino):

in our case it is negative, namely the function is shifted to the right

OpenStudy (susanna):

thanx

OpenStudy (michele_laino):

:)